can i know the value of sin and cos ?


1
Expert's answer
2021-10-06T15:52:31-0400

To find the value of sin and cos, draw a circle of radius 1 inch or 1 cm (or just one unit) like this:



Draw a line at some angle counting the angle from the positive x-direction. To find the value of "\sin\theta", draw a line from point A toward y-axis parallel to x-axis. The y-value is the value of "\sin\theta".

Draw a vertical line from point A to x-axis. The x-coordinate is the value of "\cos\theta".